OpenStudy (anonymous):

For y = x2 + 4x − 12, Determine if the parabola opens up or down. State if the vertex will be a maximum or minimum. Find the vertex. Find the x-intercepts. Describe the graph of the equation.

OpenStudy (wolf1728):

The coefficient of x² is positive so it is 'U' shaped and opens up.

OpenStudy (anonymous):

\[y=x ^{2}+4x+4-4-12\] \[y+16=\left( x+2 \right)^{2},\it is an upward parabola with vertex (-2,-16)\] It is minimum at the vertex. for x-intercepts put y=0 \[x+2=\pm 4,x=4-2=2 and x=-4-2=-6\]